OpenStudy (anonymous):

Did I get this right? Factor: a^3y+1 I got (a^y+1)(a^y-1)

ganeshie8 (ganeshie8):

(a^y+1)(a^y-1) multiply them and see if you get back the original expression

OpenStudy (anonymous):

Give me one second. :-)

OpenStudy (anonymous):

I got a^y2-1a^y-1

OpenStudy (anonymous):

I think you will get : [a(^y)^2 - 1^2] after multiplying (a^y+1)(a^y-1)

OpenStudy (anonymous):

Well, it appears that I don't know how to do it. Can someone perhaps explain. @ganeshie8

OpenStudy (anonymous):

@doddle

OpenStudy (anonymous):

Hmm it is really simple just remember this equation: a^3y+1 is a sum of cubes. a^3+b^3= (a+b) (a^2-ab+b^2) So, a^3y+1= (a^y+1)(a^2y-a+1)